76 Lexington Avenue, �1 Somerville, MA 02144

November 17, 1998

Alkermes, Inc. Human Resources 64 Sidney Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Alkermes, Inc. for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

American Biogenetic Sciences, Inc. Human Resources 801 Albany Street, Suite 300 Boston, MA 02118

Dear Sir or Madam,

I am interested in openings within American Biogenetic Sciences for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Apollo Genetics Human Resources One Kendall Square, Suite 220 Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Apollo Genetics for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Ariad Pharmaceuticals, Inc. Human Resources 26 Landsdowne Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Ariad Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Astra Research Center Boston, Inc. Human Resources 128 Sidney Street Cambridge, MA 02139-4239

Dear Sir or Madam,

I am interested in openings within the Astra Group’s Cambridge location for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

AutoImmune, Inc. Human Resources 128 Spring Street Lexington, MA 02173

Dear Sir or Madam,

I am interested in openings within AutoImmune Inc. for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Bion, Inc. Human Resources 840 Memorial Drive Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Bion, Inc. for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Boston Life Sciences, Inc. Human Resources 31 Newbury Street Boston, MA 02116

Dear Sir or Madam,

I am interested in openings within Boston Life Sciences for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Cambridge NeuroScience Human Resources One Kendall Square, Bldg. 700 Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Cambridge NeuroScience for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Circe Biomedical, Inc. Human Resources One Ledgemont Center 128 Spring Street Lexington, MA 02173

Dear Sir or Madam,

I am interested in openings within Circe Biomedical for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Cubist Pharmaceuticals, Inc. Human Resources 24 Emily Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Cubist Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Cytomatrix, Inc. Human Resources 100 Inman St., Suite 104 Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Cytomatrix for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

CytoMed, Inc. Human Resources 840 Memorial Drive Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within CytoMed for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Dyax, Inc. Human Resources One Kendall Square Building 600, Fifth Floor Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Dyax for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Ergo Science, Inc. 33 Third Ave, Third Floor Charlestown, MA 02129 Attn: Mindy Gikas, Human Resources Director

Dear Ms. Gikas,

I am interested in openings within Ergo Science Inc. for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

ETEX Corporation Human Resources 38 Sidney Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within ETEX for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Genetics Institute, Inc. Human Resources 87 Cambridge Park Drive Cambridge, MA 02140

Dear Sir or Madam,

I am interested in openings within The Genetics Institute for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Genetix Pharmaceuticals, Inc. Human Resources 840 Memorial Drive Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Genetix Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Genzyme Corporation Human Resources One Kendall Square, Bldg. 1400 Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Genzyme for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Glycan Pharmeceuticals, Inc. Human Resources 100 Inman Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Glycan Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Hybridon, Inc. Human Resources 620 Memorial Drive Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Hybridon for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

IGG International, Inc. Human Resources One Kendall Square, Bldg. 300 Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within IGG International for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

ImmunoGen, Inc. Human Resources 128 Sidney Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within ImmunoGen for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Interneuron Pharmaceuticals, Inc. Human Resources 99 Hayden Avenue, Suite 340 Lexington, MA 02173

Dear Sir or Madam,

I am interested in openings within Interneuron Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

IntraImmune Therapies, Inc. Human Resources PO Box 15599 Boston, MA 02215-0011

Dear Sir or Madam,

I am interested in openings within ITI for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Kinetix Pharmaceuticals, Inc. Human Resources 200 Boston Ave, Suite 4700 Medford, MA 02155

Dear Sir or Madam,

I am interested in openings within Kinetix Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

LeukoSite, Inc. Personnel 215 First Street Cambridge, MA 02142

Dear Sir or Madam,

I am interested in openings within LeukoSite, Inc. for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Millenium Pharmaceuticals, Inc. Human Resources 640 Memorial Drive, Fifth Floor Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Millenium for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Mitotix, Inc. Human Resources One Kendall Square, Bldg. 600 Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Mitotix for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Molecular Geodesics Human Resources 201 Broadway Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Molecular Geodesics for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

NeoGenesis Drug Discovery, Inc. Human Resources 840 Memorial Drive, Third Floor Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within NeoGenesis for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Nissin Molecular Biology Institute Human Resources 20 Overland Street Boston, MA 02215

Dear Sir or Madam,

I am interested in openings within Nissin Molecular Biology Institute for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

OmniGene Bioproducts, Inc. Human Resources 763 D Concord Ave. PO Box 9002 Cambridge, MA 02139-9002

Dear Sir or Madam,

I am interested in openings within OmniGene Bioproducts for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Oncogene Science Human Resources 80 Rogers Street Cambridge, MA 02142

Dear Sir or Madam,

I am interested in openings within Oncogene Science for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Ontogeny, Inc. 45 Moulton Street Cambridge, MA 02138 Attn: Marlene A Thayer, Dir., Human Resources

Dear Ms. Thayer,

I am interested in openings within Ontogeny Inc. for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

OraVax, Inc. Human Resources 38 Sidney Street Cambridge, MA 02139-4169

Dear Sir or Madam,

I am interested in openings within OraVax for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Pangaea Pharmaceuticals, Inc. Human Resources 763 E. Concord Ave. Cambridge, MA 02138

Dear Sir or Madam,

I am interested in openings within Pangaea Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Pentose Pharmaceuticals, Inc. Human Resources 45 Moulton Street Cambridge, MA 02138

Dear Sir or Madam,

I am interested in openings within Pentose Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Peptimmune, Inc. Human Resources 840 Memorial Drive, Third Floor Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Peptimmune for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Point Therapeutics Human Resources 75 Kneeland Street Boston, MA 02111

Dear Sir or Madam,

I am interested in openings within Point Therapeutics for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Praecis Pharmaceuticals, Inc. Human Resources One Hampshire Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Praecis Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Procept, Inc. Human Resources 840 Memorial Drive Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Procept for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Proscript, Inc. Human Resources 38 Sidney Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Proscript for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Reprogenesis Human Resources 800 Huntington Ave. Boston, MA 02115

Dear Sir or Madam,

I am interested in openings within Reprogenesis for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Scriptgen Pharmaceuticals, Inc. Human Resources 200 Boston Avenue, Suite 3000 Medford, MA 02155

Dear Sir or Madam,

I am interested in openings within Scriptgen Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Therion Biologics Corporation Human Resources 76 Rogers Street Cambridge, MA 02142

Dear Sir or Madam,

I am interested in openings within Therion Biologics for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Transcend Therapeutics, Inc. Human Resources 640 Memorial Drive, 3 West Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Transcend Therapeutics for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Transkaryotic Therapies, Inc. Human Resources 195 Albany Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Transkaryotic Therapies for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Variagenics One Kendall Square, Bldg. 400 Cambridge, MA 02139 Attn: Eleonora Brown, Human Resources

Dear Ms. Brown,

I am interested in openings within Variagenics for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Vertex Pharmaceuticals, Inc. Human Resources 130 Waverly Street Cambridge, MA 02139

Dear Sir or Madam,

I am interested in openings within Vertex Pharmaceuticals for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,

November 17, 1998

Virus Research Institute Human Resources 61 Moulton Street Cambridge, MA 02138

Dear Sir or Madam,

I am interested in openings within the Virus Research Institute for the position of associate scientist. I have extensive experience in a laboratory setting in the areas of protein chemistry, cellular biology and enzyme kinetics. Most recently I have organized and run an enzyme kinetics laboratory involved in searching for novel protease inhibitors. I have enclosed my resume for your perusal. Please contact my by email at cjsmith@mit.edu or by phone at (617) 628-0406 with questions or to set up an interview. Thank you.

Sincerely,